OpenStudy (mayaal):

Express the complex number in trigonometric form. -2i 2(cos 180° + i sin 180°) 2(cos 90° + i sin 90°) 2(cos 270° + i sin 270°) 2(cos 0° + i sin 0°)

OpenStudy (anonymous):

since you have choices it should be an easy check see which one works

OpenStudy (anonymous):

\[ 2(cos 180° + i \sin 180°)=2(-1+0)=-2\] so not that one

OpenStudy (mayaal):

but what if i didn't have those choices?what would i do then?

OpenStudy (anonymous):

then you would have a bit of work to do but not in this example because this is an easy one

OpenStudy (anonymous):

to write \[2i\] as \[r(\cos(\theta)+i\sin(\theta))\] you need two numbers, \(r\) and \(\theta\)

zepdrix (zepdrix):

I can't believe how often this question shows up lol. It's like every couple of hours I see a new person posting this exact question.

OpenStudy (agreene):

2 tells you its r= 2 -i tells you its \(\theta\) = -90

OpenStudy (anonymous):

\(r\) is easy it is \(r=|-2i|\) which is pretty clearly \(2\)

OpenStudy (anonymous):

@zepdrix ikr

OpenStudy (agreene):

pretty sure the same questions show up around the same day because a lot of the home school kids are in the same section/s and having the same issues with the more difficult problems

OpenStudy (anonymous):

you also need \(\theta\) which is easy in this case as well 

OpenStudy (anonymous):

home school an oxymoron if ever there was one more like "homescam" which i had gotten in on it though...

OpenStudy (agreene):

lol yeah, I think it would have been bad if I had gotten in on it though, probably wouldn't be where I am today if I didn't actually >>learn<< things

OpenStudy (mayaal):

haha.thnku sooooooo much guys!